[Brahms-l] Shift report 20040330 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 30 2004 - 23:53:23 EST
------------------------------------------------------------
Date: Tue Mar 30 23:53:21 2004, Shift: 16:00-24:00
Supervisor: R. Debbe, with: 
Log book#15 pages 103 - 103

Shift Summary:
------------------------------------------------------------
Two good stores
Moved MRS spectrometer to 90 and collected data with D5 at 2Kgauss Pol B
FS at 4 degrees
runs 11455 11459 11460 11466 and 11467 (still running)

DAQ had time-outs that I managed to fix by clearing alarms on VME crates (may be just a coincidence) But the next run I took was not written to HPSS (11466) hopefully somebody can move it there.

No detector problems
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 30 23:54:03 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 30 2004 - 23:54:23 EST